Session: May 2001 (session 2)
Wavelength: 3.6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (orfei at Thu May 31 14:55:49 METDST 2001) Success
No problems to comment

Torun (K.Borkowski at Thu May 31 14:56:54 METDST 2001) Success
No known problems.

Effelsberg (A. Kraus at Thu May 31 17:03:07 METDST 2001) Success

Noto (P. Leto at Sun Jun 3 18:46:45 METDST 2001) Success

Noto (P .Leto at Sun Jun 3 18:50:26 METDST 2001) Success

Westerbork (Tony Foley at Mon Jun 4 11:42:46 METDST 2001) Minor failures
Heavy rain with some lightning. FS was stuck waiting for a response from phasecal unit 10:20-11:00. This was probably caused by nearby lighnting. Phasing the array also poor with unstable weather.

Jodrell_Bank (Alastair Gunn at Mon Jun 4 12:34:23 METDST 2001) Success
No known problems.

Onsala (John Conway at Tue Jun 5 11:05:18 METDST 2001) Success
No knowsn problems